# DSO: Aligning 3D Generators with Simulation Feedback for Physical Soundness

### v1

@@ 4 Experiments
We evaluate DSO on the task of generating physically stable 3D models under gravity and compare it to prior works that use test-time optimization of physically-based losses (Sec. 4.2). We test the ability to generate stable 3D objects while also retaining the fidelity of the 3D reconstruction (as it would be trivial to make all objects stable by making them, e.g., cubes). We discuss the effect of DSO on the generated geometry in Sec. 4.3 and DSO’s scaling behavior in Sec. 4.5. In Sec. 4.6, we demonstrate how DSO can be adapted to leverage exclusively synthetic 2D images instead of renderings of ground-truth 3D models.
@@ 4.1 Experiment Details
We apply our method DSO to fine-tune TRELLIS [101], a state-of-the-art image-to-3D generator, and measure its ability to consistently generate self-supporting 3D models before and after optimization. TRELLIS contains two rectified flow transformers: the first is tasked with generating the coarse geometry of the 3D object, and the second its fine-grained details. In our experiments, we fine-tune the linear layers of the first transformer only, as we found that stability is mostly controlled by the coarse geometry. We use LoRA [25] to reduce the number of parameters to optimize. We select TRELLIS because it is a state-of-the-art 3D generator and it is available as open source, but our method is not specific to this model.
For the training data, we first generate a large number of 3D models with TRELLIS, conditioned on Objaverse [12] renderings. We exclude objects from Objaverse with unstable ground-truth shapes and filter out low-quality ones following [101]. In addition, we include only objects categorized by GObjaverse [70] as “Human-Shape”, “Animals”, or “Daily-Used”, as these categories often feature two-legged shapes and tall, slender structures, making them more challenging to stabilize under gravity. We render $6$ images for each of the remaining $13$k objects and generate $4$ different models per image, yielding $312$k 3D models in total. We then use the MuJoCo [88] simulator to conduct physical simulations for each model, starting from an upright pose on flat ground. We use its tilting angle at the final equilibrium state to determine its stability, based on a hard cut-off of $20^{\circ}$: a model $\bm{x}_{0}$ is considered stable (i.e., $o(\bm{x}_{0})=1$) if its tilting angle is below $20^{\circ}$ and unstable otherwise. During training, we sample models for an image prompt uniformly at random.
@@
We use AdamW [49] to fine-tune the base model using LoRA [25] (rank $64$) with batch size $48$ on $4$ NVIDIA A100 GPUs. We train two separate models, optimizing them using $\mathcal{L}_{\text{DRO}}$ (Eq. 5) for $4,000$ steps and using $\mathcal{L}_{\text{DPO}}$ (Eq. 7) for $8,000$ steps, respectively. The $\beta$ in Eq. 7 is set to $500$. More training details can be found in Appendix B.
@@
We evaluate on the dataset as in [21], which consists of $100$ Objaverse [12] objects from plants, animals, and characters. We exclude the $35$ objects whose ground-truth shape is not self-supporting and render $12$ images for each of the remaining objects, resulting in a final set of $65$ objects and $780$ images. These objects are removed from our training set.
@@
[…]
@@
In addition to our base model TRELLIS [101], we consider two baseline methods designed to generate self-supporting 3D objects: Atlas3D [7] and PhysComp [21]. Atlas3D is a text-to-3D framework that combines score distillation sampling [68] with physically based loss terms, primarily the magnitude of the object orientation change at equilibrium, computed via differentiable simulation. PhysComp takes a (volumetric) tetrahedral mesh as input and applies test-time optimization to improve its physical soundness, including its stability under gravity. This is achieved by encouraging the projection of the center of mass to be within the convex hull of the contact points. For [7] and [21], we use their official implementation. For the text-conditioned Atlas3D, we prompt it using captions of our multi-view renderings, obtained with GPT-4V [64]. We generate one asset per object in the evaluation set. For PhysComp, we task it to optimize the 3D models generated by TRELLIS. Since the optimization on our hardware ($24$-core CPU with $668$ GiB RAM in total) takes significantly longer (on average $15$ minutes) than the $80$ seconds reported by the authors, we only run it on an $11$-object subset whose renderings lead TRELLIS to generate unstable 3D models, amounting to $11\times 12=132$ runs. As the optimization time varies dramatically with mesh complexity, we set a strict time budget of $30$ minutes per run.
@@ 4.2 Results and Analysis
[…]
@@
TRELLIS + DSO w/ $\mathcal{L}_{\text{DPO}}$ $\underline{95.1}$ $\underline{5.42}^{\circ}$ 0.04800.0480{0.0480}0.0480 73.6273.62{73.62}73.62
@@ 4.4 Comparison with Supervised Fine-tuning
To further assess the effectiveness of DSO, we also compare it with supervised fine-tuning (SFT) in Tab. 2. For SFT, we fine-tune TRELLIS on the stable subset of our constructed dataset (i.e., $\{\bm{x}_{0}\in\mathcal{X}|o(\bm{x}_{0})=1\}$, consisting of $72$k objects out of the $312$k generated in total), using the rectified flow objective [1, 42, 46] with the same hyperparameter configuration as our main training runs for 8,000 steps. While SFT yields better geometry, its samples are less physically sound. This suggests the model prioritizes geometry over physical plausibility, making fine-tuning 3D generators solely on physically stable objects less effective for aligning physical soundness. In contrast, by exposing the model to both stable and unstable objects, DSO encourages the model to better focus on physical properties.
@@ 4.5 Scaling Behaviors
In LABEL:fig:scaling_data, we analyze the impact of training data size on model performance. We train $6$ models with identical hyperparameters as our main training run, progressively reducing the amount of data exposed to each model. The smallest dataset used is only $\frac{1}{64}$ of the full dataset, constructed as described in Sec. 4.1. While training on extremely small datasets leads to model collapse, we find that using just $\frac{1}{16}$ of the full dataset (equivalent to $19.2$k synthetic 3D models with simulation feedback) already produces results comparable to our main training run. This suggests that aligning state-of-the-art 3D generators with physical soundness requires only a modest amount of preference data. This is promising for aligning other physical properties, such as 3D scene decomposition [105, 104, 61] and part articulation [37, 38, 50], for which obtaining positive samples may be more challenging due to their rarity.
@@ 4.6 DSO without Real Data
Our training objective does not rely on ground-truth 3D data for supervision. Nevertheless, in our main experiments presented in Sec. 4.2, we used Objaverse renderings as prompts to construct a preference dataset. Here, we show that access to Objaverse models is not necessary. We substitute the renderings with object-centric synthetic images to condition the base model TRELLIS to generate 3D models. We then evaluate the physical stability of these generated models using simulation feedback, assigning a binary preference label which we use for DSO fine-tuning. In more detail, we task GPT-4 [64] to generate 1,000 diverse prompts of detailed object descriptions and use them to prompt FLUX [33], an open-source text-to-image model, to generate synthetic images. We then obtain in total $64$k generated 3D assets, on which we conduct physical simulation as detailed in Sec. 4.1. The performance of the model trained on this dataset is reported in Tab. 3. Despite the larger domain gap, the fine-tuned model generalizes well to the evaluation images and is more likely than the base model TRELLIS to generate stable assets under gravity. Without relying on physically sound 3D datasets, this self-improving scheme can be extended to enhance other physical attributes that can be evaluated via simulation for 3D generators.
@@ 5 Conclusion
We presented DSO, a new framework to generate physically sound 3D objects by leveraging feedback from a physics simulator. We use a dataset of 3D objects labeled with stability scores obtained from the simulator, potentially starting from entirely synthetic images. We fine-tune the base generator using objective DPO or DRO, the latter of which we introduced. The resulting feed-forward generator is much faster and more reliable in producing stable objects compared to test-time optimization methods.
@@ Acknowledgments.
This work is in part supported by a Toshiba Research Studentship, EPSRC SYN3D EP/Z001811/1, and ERC-CoG UNION 101001212. We thank Minghao Guo and Bohan Wang for providing us with the evaluation set in their work [21], and Mariem Mezghanni for insightful discussions during the early stages of this project. We also thank Zeren Jiang, Minghao Chen, Jinghao Zhou and Gabrijel Boduljak for helpful suggestions.
@@ Appendix C Additional Evaluation Details
Appendix D Additional Results on In-the-Wild Images
@@
Appendix E Limitations and Future Work

### v2

@@ 4 Experiments
We evaluate DSO on the task of generating physically stable 3D models under gravity and compare it to prior works that use test-time optimization of physically-based losses (Sec. 4.2). We assess the ability to generate stable 3D objects while retaining the fidelity of the 3D reconstruction (as it would be trivial to make all objects stable by making them, e.g., cubes). We discuss the effect of DSO on the generated geometry in Sec. 4.3 and DSO’s scaling behavior in Sec. 4.5. In Sec. 4.6, we demonstrate how DSO can be adapted to leverage exclusively synthetic 2D images instead of renderings of ground-truth 3D models.
@@ 4.1 Experiment Details
We apply DSO to fine-tune TRELLIS [101], a state-of-the-art image-to-3D generator, and measure its ability to consistently generate self-supporting 3D models before and after optimization. TRELLIS contains two rectified flow transformers: the first generates the coarse geometry of the 3D object, and the second refines its fine-grained details. In our experiments, we fine-tune only the linear layers of the first transformer, as stability is primarily controlled by the coarse geometry. We use LoRA [25] to reduce the number of parameters to optimize. We select TRELLIS because it is a state-of-the-art 3D generator and is available as open source, but our method is not specific to this model.
For the training data, we first generate a large number of 3D models with TRELLIS, conditioned on Objaverse [12] renderings. We exclude objects from Objaverse with unstable ground-truth shapes and filter out low-quality ones following [101]. Additionally, we include only objects categorized by GObjaverse [70] as “Human-Shape”, “Animals”, or “Daily-Used”, as these categories often feature two-legged shapes and tall, slender structures, making them more challenging to stabilize under gravity. We render $6$ images for each of the remaining 1313k objects and generate $4$ different models per image, yielding 312312k 3D models in total. We then use the MuJoCo [88] simulator to conduct physical simulations for each model, starting from an upright pose on flat ground. We use the tilting angle at the final equilibrium state to determine stability, based on a hard cut-off of $20^{\circ}$: a model $\bm{x}_{0}$ is considered stable (i.e., $o(\bm{x}_{0})=1$) if its tilting angle is below $20^{\circ}$ and unstable otherwise. During training, we sample models for an image prompt uniformly at random.
@@
We use AdamW [49] to fine-tune the base model using LoRA [25] (rank $64$) with a batch size of $48$ on 44 NVIDIA A100 GPUs. We train two separate models, optimizing them using $\mathcal{L}_{\text{DRO}}$ (Eq. 5) for $4,000$ steps and using $\mathcal{L}_{\text{DPO}}$ (Eq. 7) for $8,000$ steps, respectively. The $\beta$ in Eq. 7 is set to $500$. More details can be found in Appendix B.
@@
We evaluate on the dataset from [21], which consists of $100$ Objaverse [12] objects from plants, animals, and characters. We exclude the $35$ objects whose ground-truth shape is not self-supporting and render $12$ images for each of the remaining objects, resulting in a final set of $65$ objects and $780$ images. These objects are removed from our training set.
@@
For quantitative results, we report the following stability measures: $\%$ Output counts the frequency of successfully outputting a 3D object, regardless of its stability; $\%$ Stable counts the percentage of stable assets among those generated; Rotation angle (Rot. in short) measures the average tilting angle of generated objects at their equilibrium state. In addition, to evaluate the mesh geometry, we report Chamfer Distance (CD) and F-Score (with threshold $0.05$ [55, 94, 44]). Following common practices [55, 94, 44], we scale the meshes to fit within the unit cube and align the generated meshes optimally with the ground truths using Iterated Closest Point (ICP) before computing CD and F-Score.
@@
In addition to our base model TRELLIS [101], we consider two baseline methods designed to generate self-supporting 3D objects: Atlas3D [7] and PhysComp [21]. Atlas3D is a text-to-3D framework that combines score distillation sampling [68] with physically-based loss terms, primarily the magnitude of the object orientation change at equilibrium, computed via differentiable simulation. PhysComp takes a (volumetric) tetrahedral mesh as input and applies test-time optimization to improve its physical soundness, including its stability under gravity. This is achieved by encouraging the projection of the center of mass to be within the convex hull of the contact points. For [7] and [21], we use their official implementations. For the text-conditioned Atlas3D, we prompt it using captions of our multi-view renderings, obtained with GPT-4V [64]. We generate one asset per object in the evaluation set. For PhysComp, we task it to optimize the 3D models generated by TRELLIS. Since the optimization on our hardware ($24$-core CPU with $668$ GiB RAM in total) takes significantly longer (on average $15$ minutes) than the $80$ seconds reported by the authors, we only run it on an $11$-object subset whose renderings lead TRELLIS to generate unstable 3D models, amounting to $11\times 12=132$ runs. As the optimization time varies dramatically with mesh complexity, we set a strict time budget of $30$ minutes per run.
@@ 4.2 Results and Analysis
We note that: (1) Differentiable simulation often suffers from numerical issues, as reflected by the lower $\%$ Output of [7] and [21], due to the need for differentiable ODE solving. DSO circumvents this requirement by framing physical soundness optimization as a reward learning task (Sec. 3.2) and augmenting 3D models with simulation feedback before training. (2) Unlike visual quality, physical stability demands high accuracy, especially in the contact region. While existing efforts to align vision generators [90, 16, 43, 114, 69] focus on enhancing visual quality, we show that alignment can also substantially improve accuracy-sensitive metrics. (3) For our task, $\mathcal{L}_{\text{DRO}}$ proves to be a more effective objective than $\mathcal{L}_{\text{DPO}}$ (Tab. 1) and could also be beneficial in other diffusion alignment settings, especially when access to pairwise preference data is limited.
@@
TRELLIS + DSO w/ $\mathcal{L}_{\text{DPO}}$ $\underline{95.1}$ $\underline{5.42}^{\circ}$ 0.0480{0.0480} 73.62{73.62}
@@ 4.4 Comparison with Supervised Fine-tuning
To further assess the effectiveness of DSO, we also compare it with supervised fine-tuning (SFT) in Tab. 2. For SFT, we fine-tune TRELLIS on the stable subset of our constructed dataset (i.e., $\{\bm{x}_{0}\in\mathcal{X}|o(\bm{x}_{0})=1\}$, consisting of 7272k objects out of the 312312k generated in total), using the rectified flow objective [1, 42, 46] with the same hyperparameter configuration as our main training runs for 8,000 steps. While SFT yields better geometry, its samples are less physically sound. This suggests that the model prioritizes geometry over physical plausibility, making fine-tuning 3D generators solely on physically stable objects less effective for aligning physical soundness. In contrast, by exposing the model to both stable and unstable objects, DSO encourages the model to better focus on physical properties.
@@ 4.5 Scaling Behaviors
In LABEL:fig:scaling_data, we analyze the impact of training data size on model performance. We train $6$ models with identical hyperparameters as our main training run, progressively reducing the amount of data exposed to each model. The smallest dataset used is only $\frac{1}{64}$ of the full dataset, constructed as described in Sec. 4.1. While training on extremely small datasets leads to model collapse, we find that using just $\frac{1}{16}$ of the full dataset (equivalent to 19.219.2k synthetic 3D models with simulation feedback) already produces results comparable to our main training run. This suggests that aligning state-of-the-art 3D generators with physical soundness requires only a modest amount of preference data. This is promising for aligning other physical properties, such as 3D scene decomposition [105, 104, 61] and part articulation [37, 38, 50], for which obtaining positive samples may be more challenging due to their rarity.
@@ 4.6 DSO without Real Data
Our training objective does not rely on ground-truth 3D data for supervision. Nevertheless, in our main experiments presented in Sec. 4.2, we used Objaverse renderings as prompts to construct a preference dataset. Here, we show that access to Objaverse models is not necessary. We substitute the renderings with object-centric synthetic images to condition the base model TRELLIS to generate 3D models. We then evaluate the physical stability of these generated models using simulation feedback, assigning a binary preference label, which we use for DSO fine-tuning. In more detail, we task GPT-4 [64] to generate 1,000 diverse prompts of detailed object descriptions and use them to prompt FLUX [33], an open-source text-to-image model, to generate synthetic images. We then obtain a total of 6464k generated 3D assets, on which we conduct physical simulation as detailed in Sec. 4.1. The performance of the model trained on this dataset is reported in Tab. 3. Despite the larger domain gap, the fine-tuned model generalizes well to the evaluation images and is more likely than the base model TRELLIS to generate stable assets under gravity.
@@ 5 Conclusion
We presented DSO, a novel framework for generating physically sound 3D objects by leveraging feedback from a physics simulator. Our approach utilizes a dataset of 3D objects labeled with stability scores obtained from the simulator, potentially starting from entirely synthetic images. We fine-tune the base generator using the DPO or DRO objectives, the latter of which we introduced. The resulting feed-forward generator is significantly faster and more reliable at producing stable objects compared to test-time optimization methods.
@@ Acknowledgments.
This work is supported by a Toshiba Research Studentship, EPSRC SYN3D EP/Z001811/1, and ERC-CoG UNION 101001212. We thank Minghao Guo and Bohan Wang for providing us with the evaluation set in their work [21], and Mariem Mezghanni for insightful discussions during the early stages of this project. We also thank Zeren Jiang, Minghao Chen, Jinghao Zhou, and Gabrijel Boduljak for helpful suggestions.
@@ Appendix C Additional Evaluation Details
Appendix D Additional Results
D.1 Additional Evaluation Results
To demonstrate that the enhanced physical soundness achieved through DSO is not limited to a specific simulation environment, we report the evaluation results in Isaac Gym [53] and under perturbations in Tab. 6. For the evaluation under perturbations, we choose $4$ maximum perturbation angles $\theta_{\max}$ and perform $100$ simulation runs with each $\theta_{\max}$ where the generated 3D models are initially rotated by a random angle $\theta\in(-\theta_{\max},\theta_{\max})$, following Atlas3D [7]. We then report the average stability rate of the $100$ runs. In Tab. 6, TRELLIS post-trained with only MuJoCo feedback via DSO outperforms all baselines under all simulation settings, showing that the improved physical soundness generalizes well to different simulation environments.
Method MuJoCo Isaac Gym
w/o perturbation $\theta_{\max}=0.01$ $\theta_{\max}=0.02$ $\theta_{\max}=0.04$ $\theta_{\max}=0.08$ w/o perturbation
Full evaluation set ($65$ objects)
TRELLIS [101] $85.1$ $84.8$ $84.2$ $82.5$ $77.2$ $97.3$
Atlas3D [7] $69.4$ $70.3$ $70.2$ $66.3$ $61.8$ $88.7$
TRELLIS + DSO (w/ $\mathcal{L}_{\text{DPO}}$) $\underline{95.1}$ $\underline{94.8}$ $\underline{94.1}$ $\underline{92.6}$ $\underline{88.0}$ $\underline{99.3}$
TRELLIS + DSO (w/ $\mathcal{L}_{\text{DRO}}$) $\mathbf{99.0}$ $\mathbf{98.8}$ $\mathbf{98.6}$ $\mathbf{97.2}$ $\mathbf{93.7}$ $\mathbf{99.6}$
Partial evaluation set ($11$ unstable objects)
TRELLIS [101] $54.5$ $54.0$ $53.8$ $48.5$ $41.5$ $93.9$
TRELLIS + PhysComp [21] $80.3$ $76.9$ $76.1$ $72.6$ $\underline{67.7}$ $83.9$
TRELLIS + DSO (w/ $\mathcal{L}_{\text{DPO}}$) $\underline{82.6}$ $\underline{82.0}$ $\underline{80.7}$ $\underline{77.5}$ $67.5$ $\underline{98.5}$
TRELLIS + DSO (w/ $\mathcal{L}_{\text{DRO}}$) $\mathbf{95.5}$ $\mathbf{95.4}$ $\mathbf{95.0}$ $\mathbf{93.9}$ $\mathbf{85.4}$ $\mathbf{100.0}$
Table 6: Results evaluated under different simulation settings.
D.2 Additional Comparison with Post-Processing Baselines
In Tab. 7, we compare DSO with a naive post-processing baseline that cuts the mesh flat just above the lowest vertex, following Atlas3D [7]. This method is less effective at stabilizing meshes and significantly degrades geometric quality, as reflected in the higher Chamfer distance (Tab. 7).
Method Enforcing flat at height $z$ DSO (Ours)
$z=0.05$ $z=0.1$ $z=0.15$ $z=0.2$
$\%$ Stable $94.2$ $90.5$ $93.2$ $\underline{95.8}$ $\mathbf{99.0}$
Chamfer Distance $\underline{0.0502}$ $0.0537$ $0.0591$ $0.0662$ $\mathbf{0.0440}$
Table 7: Comparison with post-processing baselines.
@@ D.2 Additional Comparison with Post-Processing Baselines
D.3 Additional Results on In-the-Wild Images
@@
Appendix E Additional Discussions
A deeper analysis of DRO vs. DPO.
We further analyze the similarities and differences between $\mathcal{L}_{\text{DRO}}$ and $\mathcal{L}_{\text{DPO}}$. Both losses are monotonic functions of $o=\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\theta}(\bm{x}_{t}^{w},t)\|^{2}_{2}-\|\bm{\epsilon}^{w}-\bm{\epsilon}_{\text{ref}}(\bm{x}_{t}^{w},t)\|^{2}_{2}-\left(\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\theta}(\bm{x}_{t}^{l},t)\|^{2}_{2}-\|\bm{\epsilon}^{l}-\bm{\epsilon}_{\text{ref}}(\bm{x}_{t}^{l},t)\|^{2}_{2}\right)$. In Fig. 8, we plot each loss (left) and its derivative with respect to $o$ (right, log-scale). A key difference is that $\frac{d\mathcal{L}_{\text{DRO}}}{do}$ is constant, while $\frac{d\mathcal{L}_{\text{DPO}}}{do}$ decays exponentially as $o$ decreases. As a result, $o$ tends to plateau during optimization of $\mathcal{L}_{\text{DPO}}$. This leads to faster convergence with $\mathcal{L}_{\text{DRO}}$, although extended training may harm performance.
Figure 8: Plots of $\mathcal{L}_{\text{DRO}}$ and $\mathcal{L}_{\text{DPO}}$ and their derivatives.
Scaling behaviors when optimizing $\mathcal{L}_{\text{DRO}}$.
In Sec. 4.5, we analyzed how DSO scales when optimizing $\mathcal{L}_{\text{DPO}}$. Here, we present the corresponding scaling behavior for $\mathcal{L}_{\text{DRO}}$. As shown in Tab. 8, performance peaks at $4,000$ training steps, after which the geometry quality noticeably degrades—consistent with our earlier analysis. Scaling with training data follows a similar trend to that observed for $\mathcal{L}_{\text{DPO}}$ in LABEL:fig:scaling_data.
Training steps $2000$ $3000$ $4000$ $5000$
$\%$ Stable $91.5$ $96.9$ $\mathbf{99.0}$ $98.7$
Chamfer D. $0.0473$ $0.0464$ $\mathbf{0.0440}$ $\color[rgb]{1,0,0}0.0853$}
Table 8: Scaling behavior with training compute of $\mathcal{L}_{\text{DRO}}$.
Appendix F Limitations and Future Work